OpenStudy (anonymous):

is (x-2) a factor of the polynomial P(x)=3x^6-192?

OpenStudy (anonymous):

if so P(2) = 0 P(2) = 3(2)^6 - 192 = 3*64 - 192 = 0 so by factor theorem (x-2) is a factor

hero (hero):

Oh wow, look at that. I'm going to have to study factor theorem

OpenStudy (anonymous):

yes - its handy - its really a special case of the remainder theorem

hero (hero):

Now, that you've mentioned that, I think I get it